# muography; non-destructive testing; machine learning; image processing; upsampling; semantic segmentation; simulation; cosmic rays

### v1

@@ 2 Materials and Methods
Training a supervised machine learning model to upsample muography images of concrete blocks requires a large dataset of paired images. Therefore we must first create a dataset that contains different sizes, orientations and types of rebar grids and tendon ducts, as well as a variety of different sampling rates. Spherical air voids were also added to the geometry design in order to assess detection ability — as a precursor to a complimentary study to look at defects — due to their simplicity and small size.
@@
The design of the Geant4 simulations to produce a large and varied dataset of 2D muography images from 700 unique concrete block designs is outlined below in Section 2.1. Each concrete block was simulated using 100 days’ worth of muons, resulting in a total dataset size of 70,000  500$\times$500 images, each with 100 different versions corresponding to sampling rates from one day to 100 days. Simulating an equivalent sampling time of 100 days was arbitrary but intentionally overestimated for this initial study to ensure sufficient sampling for a clear ground truth; by comparison, the concrete block in Niederleithinger et al. Niederleithinger et al. (2021) was imaged for only 50 days. This data was used to train the cWGAN-GP models (outlined below in Section 2.2), using the 100-day image as the ground truth and the input being randomly selected from one of the sampling rates from 1–99, ensuring sampling rate generalisation.
@@
One challenging task was to assess the perceptual quality of the output images. While there are a wide variety of metrics that can be used to assess global image quality of outputs, such as Structural Similarity Index Measure (SSIM) and Peak Signal-to-Noise Ratio (PSNR), they do not give a detailed breakdown about how the model is affecting perception of features, such as rebar grids and tendon ducts. If we want to know how the upsampling model is affecting the perception of features within the concrete, we must find a way of separating features from one another. A solution to this problem is to train a second model to perform semantic segmentation, a machine learning technique used to classify pixels within an image and assign them to a specific class. This means that each pixel in the muographic image is assigned to either concrete, rebar, tendon duct, air void or ‘unknown’. ‘Unknown’ objects were added as the fourth and final object type such that the segmentation model could assign a class to miscellaneous objects it deemed to not fall into any of the other classes. The shape, size and materials used for ‘unknown’ objects is less constrained, hence a variety of shapes and densities were used to define them. To train a model to produce segmentation maps, a set of images detailing ‘true’ object placement of features is required. Due to the use of Geant4 Monte-Carlo simulations, ‘true’ geometry information is available to generate the segmentation ground truths. For simplicity reasons, all components of the tendon ducts were assigned to one class.
@@
Assessing the performance of segmentation models is much less ambiguous than that of an I2I task like denoising or upsampling, as each class can be split up into multiple binary masks from which to calculate metrics. One of the most popular metrics for the assessment of segmentation models is the Dice-Sørensen coefficient Dice (1945); Sørensen (1948) (hereafter referred to as Dice), which is used to gauge the similarity between two sets of data:
$Dice=\frac{2TP}{2TP+FP+FN},$ (3)
@@
where $TP$ are the number of true positive pixels, $FP$ the number of false positive pixels, and $FN$ the number of false negative pixels in a given binary image.
@@ 2.1 Simulating a Dataset
The detector system used in the Geant4 simulations was that of the muon imaging system (MIS) in place at the University of Glasgow, which is a scintillating-fibre tracking system Clarkson et al. (2014). It consists of four detection modules, each consisting of two orthogonal planes of 1024 2 mm-pitch polystyrene-based plastic scintillating fibres with polymethylmethacrylate optical cladding, coupled to Hamamatsu H12700A multi-anode photomultiplier tubes (PMTs). When a muon passes through a fibre, it imparts energy, causing photoluminescence which is detected by the PMTs. A muon track is recorded when coincident muon hits are detected in all four detector planes. The MIS has an imaging area of 1066 mm×$\times$1066 mm and was set up with a vertical spacing of 530 mm between the upstream and downstream modules, between which the concrete block samples of 1000 mm×$\times$1000 mm×$\times$200 mm were placed. Within this block, the four different object types were placed: rebar grids, tendon ducts, air voids and ‘unknowns’. The subsections immediately below outline the parameters which constrained individual object design and placement by a concrete interior randomisation algorithm. Objects’ parameters were first generated — through sampling from uniform distributions — and then placed in order of rebar–duct–void–unknown, where objects would only be placed if they fit within the boundaries of the concrete block and did not overlap other objects. In order to preserve the uniform distribution of the number of objects per volume, overlapping objects would have their parameters re-sampled until they were successfully placed or the number of placement attempts exceeded 1000.
A total of 700 concrete samples were each imaged for 100 days worth of muons, with resultant vector pairs processed using PoCA to calculate scattering angles. The resultant volume of scattering angles was subsequently voxelised into 2 mm cubes. Image slices of the X-Y plane for each voxelised geometry were saved for intervals of one day of sampling. This resulted in the creation of 70,000 unique images, of which there were 100 different versions to reflect each cumulative day of sampling.
@@ 2.1.1 Carbon-Steel Rebar Grids
The cylindrical steel rods that make up rebar grids were defined in Geant4 to have a uniform density of 7.84 $\mathrm{g}\cdot\mathrm{cm}^{-3}$, with typical chemical composition within the limits of British standard grade 500B BS 4449. The diameters of these rods were kept the same for all rods within the same grid, chosen from standard rebar sizing diameters of 8, 10, 12, 16, 20 and 25 mm. Square grid spacing (100 mm, 150 mm, 200 mm, 250 mm) between rods was also randomised, of which there were a random number of between 2–12 in the X or Y direction. Between one and four different rebar grids were positioned parallel to the X-Y plane for each concrete block configuration. The minimum of a single grid here ensures that each configuration had at least one feature.
@@ 2.1.2 Tendon Ducts
Tendon ducts are large cylindrical objects that were placed such that they span the 1000 mm volume in either the X-Z or Y-Z planes. They contain three sub-components: casing, grout and steel strands. The casing material was randomly chosen between a 0.5 mm thick galvanised carbon-steel casing (7.97 $\mathrm{g}\cdot\mathrm{cm}^{-3}$), 3 mm thick High Density Polyethylene (HDPE) with a density of 0.94 $\mathrm{g}\cdot\mathrm{cm}^{-3}$ or 3 mm thick High Density Polypropylene (HDPP) with a density of 0.90 $\mathrm{g}\cdot\mathrm{cm}^{-3}$. Casing diameter was sampled from from that of 50, 60, 70, 80, 90 and 100 mm, a value which also determined the number of high tensile steel strands (density of 7.85 $\mathrm{g}\cdot\mathrm{cm}^{-3}$) contained within the duct. The grout which filled the space inside the casing had a bulk density of 1.30 $\mathrm{g}\cdot\mathrm{cm}^{-3}$, using British Standard Grade C40/50 BS 8500-1.
@@ 2.1.3 Air Voids
Air pockets caused by improper concrete pouring are a defect that can reduce the structural integrity of reinforced concrete Sun et al. (2022). Due to the simplicity of this defect, spherical air voids were placed in the concrete in order to gauge muography’s resolving capacity, in lieu of a future study. Each volume contained 0–3 spherical air voids, with diameters ranging from 10–100 mm.
@@
Between 0–2 ‘unknowns’ were added as a control to the dataset consisting of objects we would not find in reinforced concrete to add variety and shape for model generalisation. These consisted of boxes, cylinders and spheres which did not exceed a 3D bounding box with edges limited between 35 mm and 75 mm. The densities of these objects were randomly selected from water, aluminium, iron, lead and uranium and were randomly oriented upon placement in the volume.
@@ 2.2 Pix2pix with WGAN-GP
The pix2pix model is a conditional generative adversarial network (cGAN) that utilizes a U-Net for its generator and a PatchGAN for its discriminator Isola et al. (2017). In the base pix2pix cGAN model, the generator minimizes a combined loss:
@@
where $\mathcal{L}_{\text{pixel}}$ is the pixel-wise loss function of the generator, $\lambda_{\text{pixel}}$ the weighting of the pixel-wise loss and $\mathcal{L}_{\text{adv}}$ the adversarial loss based on binary classification in the discriminator (real/fake). This binary classification provides the generator with little information on how close the distribution of generated images is to true distribution. To better quantify this difference we can use the Wasserstein distance, also known as the Earth-mover’s distance Arjovsky et al. (2017). This measures the performance of the discriminator, increasing the smoothness of gradients and importantly providing informative feedback for training the generator. Unlike binary classification, the Wasserstein distance is computed using the Kantorovich-Rubinstein duality, which restricts the discriminator to be a 1-Lipschitz function. This constraint ensures that the discriminator’s gradients are stable and meaningful, which is important for the generator to learn effectively.
The gradient penalty (GP) is the preferred solution for satisfying the Lipschitz constraint in WGANs Gulrajani et al. (2017). It directly penalises the gradients of the discriminator to ensure they remain close to 1. Specifically, the gradient penalty is computed as:
@@
where $\hat{x}$ represents interpolated samples between real and generated data and $D(\hat{x})$ is the output of the discriminator. By applying this penalty, the discriminator is constrained to satisfy the 1-Lipschitz condition, ensuring stable gradients and improving the overall training stability. This leads to smoother, more informative gradients for the generator, enabling it to produce higher quality images. In the case of WGAN-GP, the generator’s loss function, $\mathcal{L}_{\text{G}}$ is updated to include the gradient penalty:
@@
where $\mathcal{L}_{\text{adv}}$ is the adversarial loss based on the Wasserstein distance, $\mathcal{L}_{\text{pixel}}$ is the pixelwise loss and $\mathcal{L}_{\text{GP}}$ is the gradient penalty term which is weighted by $\lambda_{\text{GP}}$.
The U-Net generator architecture remains similar to pix2pix, with five encoder and decoder blocks each consisting of 4$\times$4 convolution–batch normalisation–ReLU activation. The upsampling model’s generator pixelwise loss uses the mean absolute error (MAE or L1) loss:
@@
where $N$ is the number of samples, $\hat{y}_{i}$ and $y_{i}$ are the predicted value and ground truth for the $\mathrm{i^{th}}$ sample. The segmentation model uses a custom pixelwise loss function of an evenly weighted cross-entropy loss and Dice loss (which is one minus the Dice score),
@@
which balances pixelwise prediction accuracy (via cross-entropy) and structural similarity (via Dice). Since the background class (concrete) makes up a significant proportion of all pixels, the model is biased to focus on these pixels. Therefore, the Dice loss calculation excludes this class, consisting of evenly weighted contributions of the four object classes: rebar grids, tendon ducts, air voids and ‘unknowns’. Both models were trained for 100 epochs (one epoch is a full iteration of the training set), with a batch size of 48. The upsampling model training set had each of its input images re-sampled from one of its 100 versions at the start of each epoch, in order to achieve generalisation to different sampling rates. The upsampling model generates outputs $\hat{y}$ that approximate the expected outputs $y$ of 100 days of sampling. We therefore train the segmentation model to take inputs $y$ instead of upsampled images $\hat{y}$, as we aim to assess the performance of the $\hat{y}$ outputs without introducing any bias from the upsampling process. An initial learning rate of 0.01 was used for the generator and 0.001 for the discriminator, after which a step scheduler was used to steady model convergence thus reducing the learning rates by an order of magnitude every 25 epochs.
@@ 3.1 Upsampling Model Results
The upsampling model aims not only to enhance the perceptual characteristics of low-quality muographic images but also to reduce image noise — a key issue identified in Niederleithinger et al. (2021) when comparing muography to other NDT techniques. To evaluate the quality difference between images, we use two common computer vision metrics: the Structural Similarity Index Measure (SSIM) Wang et al. (2004), which assesses perception, and the Peak Signal-to-Noise Ratio (PSNR), which is widely used in signal processing and imaging to quantify image noise. These metrics provide insights not only on the upsampling model’s performance but also on the quality of the input muography images as they are sampled over time.
Using the fact that we have different sampling rate versions for each image in our dataset, we use the aforementioned metrics to look at (a) the effect of muography sampling time with respect to image quality and (b) the performance of the upsampling model on different sampling rates. We average the SSIM and PSNR metrics of the 6900 images in the test dataset — calculated with respect to the 100-day ground truth — for both inputs and outputs of the model for each sampling rate. Results are shown in Figure 1. We see that 1-day upsampled images exhibit SSIM and PSNR scores of 0.88 and 37dB, respectively. This is equivalent of the perceptual quality of a 21-day unaltered image and the noise quality of a 31-day unaltered image. However, improvements diminish as the sampling rate increases, with both input and upsampled images for SSIM and PSNR converging at a sampling rate of 55 and 85 days respectively. Visual inspection of images before and after upsampling confirms the improvements indicated by the SSIM and PSNR metrics, even at the lowest quality images — from the 1-day sampling rate.
Figure 1: Average structural similarity index measure (SSIM) and peak signal-to-noise ratio (PSNR) metrics of the test dataset, for each sampling rate. The original dataset performance is indicated by blue circular points, where the upsampled outputs of the model are indicated by red triangular points.
Demonstrating these drastic improvements, Figure 2 compares five 1-day undersampled X-Y plane images to their respective outputs from the upsampled model, as well as their associated 100-day ground truth for reference. These results confirm that the upsampling model does an excellent job of identifying feature patterns across all sampling rates to create clearer and more refined representations of undersampled muography images.

### v2

@@ 2 Materials and Methods
Training a supervised machine learning model to upsample muography images of concrete blocks requires a large dataset of paired images. Therefore, we must first create a dataset that contains different sizes, orientations, and types of rebar grids and tendon ducts, as well as a variety of different equivalent sampling times. Spherical air voids were also added to the geometry design in order to assess detection ability—as a precursor to a complimentary study to look at defects—due to their simplicity and small size.
@@
The design of the Geant4 simulations to produce a large and varied dataset of 2D muography images from 700 unique concrete block designs is outlined below in Section 2.1. Each concrete block was simulated using 100 days worth of muons, resulting in a total dataset size of 70,000 500 $\times$ 500 images, each with 100 different versions corresponding to equivalent sampling times of 1 day to 100 days. The total image dataset therefore contained seven million images, requiring over a trillion muon events to be simulated. Simulating a maximum equivalent sampling time of 100 days was arbitrary but intentionally overestimated for this initial study to ensure sufficient sampling for a clear ground truth; by comparison, the concrete block in Niederleithinger et al. Niederleithinger et al. (2021) was imaged for only 50 days. These data were used to train the cWGAN-GP models (outlined below in Section 2.2) using the 100-day image as the ground truth and the input being randomly selected from one of the equivalent sampling times of 1–99 days, ensuring image sampling time generalization.
@@
One challenging task was to assess the perceptual quality of the output images. While there are a wide variety of metrics that can be used to assess global image quality of outputs, such as the Structural Similarity Index Measure (SSIM) and Peak Signal-to-Noise Ratio (PSNR), these metrics do not give a detailed breakdown about how the model is affecting the perception of features, such as rebar grids and tendon ducts. If we want to know how the upsampling model is affecting the perception of features within the concrete, we must find a way of separating features from one another. A solution to this problem is to train a second model to perform semantic segmentation, which is a machine learning technique used to classify pixels within an image and assign them to a specific class. This means that each pixel in the muographic image is assigned to either concrete, rebar, tendon duct, air void, or ‘unknown’. ‘Unknown’ objects were added as the fourth and final object type such that the segmentation model could assign a class to miscellaneous objects it deemed to not fall into any of the other classes. The shape, size, and materials used for ‘unknown’ objects are less constrained; hence, a variety of shapes and densities were used to define them. To train a model to produce segmentation maps, a set of images detailing ‘true’ object placement of features is required. Due to the use of Geant4 Monte Carlo simulations, ‘true’ geometry information is available to generate the segmentation ground truths. For simplicity reasons, all components of the tendon ducts were assigned to one class.
@@
Assessing the performance of segmentation models is much less ambiguous than that of an I2I task like denoising or upsampling, as each class can be split up into multiple binary masks from which to calculate metrics. One of the most popular metrics for the assessment of segmentation models is the Dice–Sørensen coefficient Dice (1945); Sørensen (1948) (hereafter referred to as Dice), which is used to gauge the similarity between two sets of data:
$\mathrm{Dice}=\frac{2TP}{2TP+FP+FN},$ (3)
@@
where $TP$ defines the number of true positive pixels, $FP$ defines the number of false positive pixels, and $FN$ defines the number of false negative pixels in a given binary image.
@@ 2.1 Simulating a Dataset
The detector system used in the simulations was that of the Muon Imaging System (MIS) in place at the University of Glasgow, which is a scintillating-fibre tracking system Clarkson et al. (2014). It consists of four detection modules, each consisting of two orthogonal planes of 1024 2 millimetre-pitch polystyrene-based plastic scintillating fibres with polymethylmethacrylate optical cladding coupled to Hamamatsu H12700A multi-anode Photomultiplier Tubes (PMTs). When a muon passes through a fibre, it imparts energy, causing photoluminescence which is detected by the PMTs. A muon track was recorded when coincident muon hits were detected in all four detector planes. The MIS has an imaging area of 1066 mm $\times$ 1066 mm and was set up with a vertical spacing of 530 mm between the upstream and downstream modules, between which the concrete block samples of 1000 mm $\times$ 1000 mm $\times$ 200 mm were placed. Within this block, the four different object types were placed: rebar grids, tendon ducts, air voids, and ‘unknowns’. The subsections immediately below outline the parameters which constrained individual object design and placement by a concrete interior randomisation algorithm. Objects’ parameters were first generated—through sampling from uniform distributions—and then placed in order of rebar–duct–void–unknown, where objects would only be placed if they fit within the boundaries of the concrete block and did not overlap other objects. In order to preserve the uniform distribution of the number of objects per volume, overlapping objects had their parameters re-sampled until they were successfully placed or the number of placement attempts exceeded 1000.
This setup was modelled in Geant4, where Ecomug was used as the cosmic-ray event generator. A generating plane was placed directly above the detector from which only muons were generated at randomised positions. The event generator was configured with its default settings, where the angular distribution was defined with a zenith angle $\theta$ spanning from 00 to $\pi/2$ and an azimuthal angle $\phi$ ranging from 00 to $2\pi$. The charge ratio of positive to negative muons was set to 1.28, which followed a momentum distribution ranging from 10 MeV to 1 TeV.
A total of 700 concrete samples were each imaged for 100 days worth of muons, with resultant vector pairs processed using the PoCA to calculate scattering angles. The resultant volume of scattering angles was subsequently voxelised into 2 mm cubes. Image slices of the X–Y plane for each voxelised geometry were saved for intervals of one day of sampling. This resulted in the creation of 70,000 unique images, of which there were 100 different versions to reflect each cumulative day of sampling.
@@ 2.1.1 Carbon-Steel Rebar Grids
The cylindrical steel rods that make up rebar grids were defined in Geant4 to have a uniform density of 7.84 $\mathrm{g}\cdot\mathrm{cm}^{-3}$, with typical chemical composition within the limits of British standard grade 500B BS 4449. The diameters of these rods were kept the same for all rods within the same grid, which were chosen from standard rebar sizing diameters of 8, 10, 12, 16, 20, and 25 mm. Square grid spacing (100 mm, 150 mm, 200 mm, 250 mm) between rods was also randomised, of which there were a random number of between 2 and 12 in the X or Y direction. Between one and four different rebar grids were positioned parallel to the X–Y plane for each concrete block configuration. The minimum of a single grid here ensured that each configuration had at least one feature.
@@ 2.1.2 Tendon Ducts
Tendon ducts consisted of large cylindrical objects that were placed such that they spanned the 1000 mm volume in either the X–Z or Y–Z planes. They contained three subcomponents: casing, grout, and steel strands. The casing material was randomly chosen between a 0.5 mm thick galvanised carbon-steel casing (7.97 $\mathrm{g}\cdot\mathrm{cm}^{-3}$), 3 mm thick High Density Polyethylene (HDPE) with a density of 0.94 $\mathrm{g}\cdot\mathrm{cm}^{-3}$, or 3 mm thick High Density Polypropylene (HDPP) with a density of 0.90 $\mathrm{g}\cdot\mathrm{cm}^{-3}$. Casing diameter was sampled from from that of 50, 60, 70, 80, 90, and 100 mm, a value which also determined the number of high tensile steel strands (density of 7.85 $\mathrm{g}\cdot\mathrm{cm}^{-3}$) contained within the duct. The grout which filled the space inside the casing had a bulk density of 1.30 $\mathrm{g}\cdot\mathrm{cm}^{-3}$ using British Standard Grade C40/50 BS 8500-1.
@@ 2.1.3 Air Voids
Air pockets caused by improper concrete pouring are a defect that can reduce the structural integrity of reinforced concrete Sun et al. (2022). Due to the simplicity of this defect, spherical air voids were placed in the concrete in order to gauge muography’s resolving capacity in lieu of a future study. Each volume contained 0–3 spherical air voids, with diameters ranging from 10 to 100 mm.
@@
Between 0 and 2 ‘unknowns’ were added as a control to the dataset consisting of objects we would not find in reinforced concrete to add variety and shape for model generalization. These consisted of boxes, cylinders, and spheres which did not exceed a 3D bounding box with edges limited between 35 mm and 75 mm. The densities of these objects were randomly selected from water, aluminium, iron, lead, and uranium and were randomly orientated upon placement in the volume.
@@ 2.2 Pix2pix with WGAN-GP
The pix2pix model is a conditional Generative Adversarial Network (cGAN) that utilises a U-Net for its generator and a PatchGAN for its discriminator Isola et al. (2017). In the base pix2pix cGAN model, the generator minimises a combined loss:
@@
where $\mathcal{L}_{\text{pixel}}$ is the pixelwise loss function of the generator, $\lambda_{\text{pixel}}$ is the weighting of the pixelwise loss, and $\mathcal{L}_{\text{adv}}$ is the adversarial loss based on binary classification in the discriminator (real/fake). This binary classification provides the generator with little information on how close the distribution of generated images is to true distribution. To better quantify this difference, we can use the Wasserstein distance, also known as the Earth mover’s distance Arjovsky et al. (2017). This measures the performance of the discriminator, increasing the smoothness of gradients and importantly providing informative feedback for training the generator. Unlike binary classification, the Wasserstein distance is computed using the Kantorovich–Rubinstein duality, which restricts the discriminator to be a 1-Lipschitz function. This constraint ensures that the discriminator’s gradients are stable and meaningful, which is important for the generator to learn effectively.
The gradient penalty is the preferred solution for satisfying the Lipschitz constraint in WGANs Gulrajani et al. (2017). It directly penalises the gradients of the discriminator to ensure they remain close to 1. Specifically, the gradient penalty is computed as follows:
@@
where $\hat{x}$ represents interpolated samples between real and generated data, and $D(\hat{x})$ is the output of the discriminator. By applying this penalty, the discriminator is constrained to satisfy the 1-Lipschitz condition, ensuring stable gradients and improving the overall training stability. This leads to smoother, more informative gradients for the generator, enabling it to produce higher quality images. In the case of WGAN-GP, the generator’s loss function, $\mathcal{L}_{\text{G}}$, is updated to include the gradient penalty:
@@
where $\mathcal{L}_{\text{adv}}$ is the adversarial loss based on the Wasserstein distance, $\mathcal{L}_{\text{pixel}}$ is the pixelwise loss, and $\mathcal{L}_{\text{GP}}$ is the gradient penalty term which is weighted by $\lambda_{\text{GP}}$.
The U-Net generator architecture remains similar to pix2pix, with five encoder and decoder blocks each consisting of 4 $\times$ 4 convolution—batch normalization—ReLU activation. The upsampling model’s generator pixelwise loss uses the mean absolute error (MAE or L1) loss:
@@
where $N$ is the number of samples, and $\hat{y}_{i}$ and $y_{i}$ are the predicted value and ground truth for the ith sample. The segmentation model uses a custom pixelwise loss function of an evenly weighted cross-entropy loss and Dice loss (which is one minus the Dice score),
@@
which balances pixelwise prediction accuracy (via cross-entropy) and structural similarity (via Dice). Since the background class (concrete) makes up a significant proportion of all pixels, the model is biased to focus on these pixels. Therefore, the Dice loss calculation excludes this class, consisting of evenly weighted contributions of the four object classes: rebar grids, tendon ducts, air voids, and ‘unknowns’. Both models were trained for 100 epochs (one epoch is a full iteration of the training set), with a batch size of 48. The upsampling model training set had each of its input images re-sampled from one of its 100 versions at the start of each epoch in order to achieve generalization to different sampling times. The upsampling model generates outputs $\hat{y}$ that approximate the expected outputs $y$ of 100 days of sampling. We therefore trained the segmentation model to take inputs $y$ instead of upsampled images $\hat{y}$, as we aimed to assess the performance of the $\hat{y}$ outputs without introducing any bias from the upsampling process. An initial learning rate of 0.01 was used for the generator and 0.001 for the discriminator, after which a step scheduler was used to steady model convergence, thus reducing the learning rates by an order of magnitude every 25 epochs.
@@ 3.1 Upsampling Model Results
The upsampling model aims not only to enhance the perceptual characteristics of low-quality muographic images but also to reduce image noise—a key issue identified in Niederleithinger et al. (2021) when comparing muography to other NDT techniques. To evaluate the quality difference between images, we used two common computer vision metrics: the Structural Similarity Index Measure (SSIM) Wang et al. (2004), which assesses perception, and the Peak Signal-to-Noise Ratio (PSNR), which is widely used in signal processing and imaging to quantify image noise. These metrics provide insights not only on the upsampling model’s performance but also on the quality of the input muography images as they are sampled over time.
Utilizing the different equivalent sampling time versions for each image in our dataset, we used the aforementioned metrics to look at (a) the effect of the muography sampling time with respect to the image quality and (b) the performance of the upsampling model on different sampling times. We averaged the SSIM and PSNR metrics of the 6900 images in the test dataset—calculated with respect to the 100-day ground truth—for both the inputs and outputs of the model at each sampling time. The results are shown in Figure 1. We see that the 1-day upsampled images exhibited SSIM and PSNR scores of 0.88 and 37dB, respectively. This is equivalent to the perceptual quality of a 21-day unaltered image and the noise quality of a 31-day unaltered image. However, improvements diminished as the sampling time increased, with both the input and upsampled images for the SSIM and PSNR converging at an equivalent sampling time of 55 and 85 days, respectively. Visual inspection of the images before and after upsampling confirmed the improvements indicated by the SSIM and PSNR metrics, even at the lowest quality images—from the 1-day equivalent sampling time.
Figure 1: Average Structural Similarity Index Measure (SSIM) and Peak Signal-to-Noise Ratio (PSNR) metrics of the test dataset as a function of equivalent sampling time. The original dataset performance is indicated by blue circular points, where the upsampled outputs of the model are indicated by red triangular points.
Demonstrating these drastic improvements, Figure 2 compares five 1-day undersampled X–Y plane images to their respective outputs from the upsampled model, as well as their associated 100-day ground truth values for reference.
These results confirm that the upsampling model does an excellent job of identifying feature patterns across all sampling times to create clearer and more refined representations of undersampled muography images.